Help to solve
-8x=18

Answer by jim_thompson5910(35256)

Divide both sides by -8 to get

-8x=18

x=18/(-8)

x = -9/4 ... don't forget to reduce

Final answer: x = -9/4
--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------